Second question


2.consider a lottery with two equally likely outcomes ,$25 and $144,with associated utility function of individuals A,Ua(m)=root of 100m and individual B,Ub(m)=5m square


A.determine expected utility and utility of expect ed value of the lottery for each individual ?


B.define and compare certainty equivalent of individual A and B ?


C.compare risk premium of individual A and B ? why risk premium of individual A is higher ?


D.Compare and interpret the absolute risk aversion ,rA(m) of the two persons ?


3).consider a farmer in a rural village and a business man in an urban area that adopt various risk reduction strategies in Ethiopia .Assume also that most farmers in Ethiopia do not insure for crop and livestock failure and businessmen also hardly insured against property loss.


A) suggest at least two strategies adopted to reduce risk by each agent ?


B) explain concisely how each strategy can reduce risk by agents ?



1
Expert's answer
2022-01-26T07:03:52-0500

A=$25

B=$144

"U(A)=\sqrt{100m}"

"U(B)=5m^{2}"

a)"Expected utility=1\/2\u00d725^{1\/2} +1\/2\u00d7144^{1\/2}"

"1\/2 \u00d75 + 1\/2 \u00d712=2.5+6"

"=8.5"

"Utility of expected value"

"For individual A, 1/2 ×\sqrt{100}=5m"

"For individual B,1\/2\u00d75=2.5m"

b) by definition, utility of the certain equivalent must be equal to the expected utility.

c)"Expected utility=utility of A- risk premium of A"

"8.5=5-R.P of A"

"R.P of A=-3.5"

"8.5=2.5-R.P of B"

"R.P of B =-6"

The risk premium for A is higher owing to the certainty equivalent and compensation of returns.

d)the certainty equivalent of A is higher hence the risk aversion is low compared to B.

3) a) Risk transfer, risk reduction

)Risk Transfer: Risk transfer involves moving the risk to another third party or entity. Risk transfers can be outsourced

moved to an insurance agency, or given to a new entity as is what happens when leasing property. Risk transfers don’t always result in lower costs. Instead, a risk transfer is the best option when it can be used to reduce future damage. So, insurance can cost money, but it may end up being more cost-effective than having the risk occur and being solely responsible for reparations.


Risk reduction- Businesses can assign a level at which risk is acceptable, which is called the residual risk level. Risk reduction is the most common strategy


because there is usually a way to at least reduce risk. It involves

taking countermeasures to decrease the impact of consequences. For example, one form of risk reduction is risk transfer, like that of buying insurance. This can help the business in the case of losses.